OpenStudy (anonymous):

Please help Simplify the radical expression, use absolute value symbols when needed: ^5square root sign x^5 y^10

OpenStudy (anonymous):

\[\sqrt[5]{x^5} y^10\]

OpenStudy (anonymous):

\[\sqrt[5]{x^5y^{10}}\] \[\sqrt[5]{x^5}=x\] and \[\sqrt[5]{y^{10}}=\sqrt[5]{y^5\times y^5}=\sqrt[5]{y^5}\times \sqrt[5]{y^5}=y\times y = y^2\]

OpenStudy (anonymous):

for a final answer of \[\sqrt[5]{x^5y^{10}}=xy^2\]

OpenStudy (anonymous):

Thank you very much! :)